# Complexity-aware fine-tuning

### v2

@@ Abstract
General-purpose Large Language Models (LLMs) are frequently fine-tuned through supervised fine-tuning (SFT) to enhance performance in specific domains. Better results can be achieved by distilling the chain-of-thought of a larger model at the cost of numerous expensive calls and a much greater amount of data. We propose a novel blueprint for efficient fine-tuning that uses reasoning only for complex data identified by entropy. Specifically, across two small open models ($\approx 3B$) we split the training data into complexity categories by a single token answer entropy (ROC AUC $0.73$), fine-tune large language models (LLMs) via SFT and distillation, and show that our pipeline significantly outperforms the standard SFT approach ($0.58$ vs $0.45$ average accuracy) and outperforms the distillation approach ($0.58$ vs $0.56$ average accuracy) while using $81\%$ less data. We publish our code [^1] https://github.com/LabARSS/complexity-aware-fine-tuning/tree/eacl2026 and data [^2] https://github.com/LabARSS/complexity-aware-fine-tuning/tree/eacl2026?tab=readme-ov-file#data to facilitate further research in this direction.
@@ 1 Introduction
To address this gap, we propose a fully automated pipeline, presented in Figure 1, that focus on both accuracy and efficiency of the fine-tuning procedure. It consists of two steps: (1) a split of the data for fine-tuning by the complexity of the task to regular and hard questions and (2) training a model via supervised fine-tuning (SFT) followed by another round of SFT that utilizes distillation on a chain-of-thought from a larger model Li et al. (2023); Hsieh et al. (2023). We confirm the effectiveness of the framework by fine-tuning two open models: Qwen2.5-3B Yang et al. (2024a) and Phi-4-Mini Microsoft et al. (2025), on the multiple choice question answering dataset MMLU-Pro Wang et al. (2024).
@@
  • A training pipeline to fine-tune LLMs via SFT and distillation for regular and hard data correspondingly. Our procedure outperforms the standard SFT approach with accuracies $0.52$ and $0.64$ over two LLMs for MMLU-Pro c.t. $0.39$ and $0.51$ for SFT and $0.50$ and $0.63$ for distillation baselines.
@@ 3.1 Training pipeline
Data aggregation.
To aggregate the data into groups by complexity (regular, hard), we evenly divide the dataset into three parts, ordering the entries by entropy of the response. A group with lower entropy values is categorized as easy and with the higher values as hard, as Figure 2 depicts.
Figure 2: Data aggregation: we can further split the data to various complexity chunks
Fine-tuning.
We propose to apply different fine-tuning strategies according to the complexity of the data.
For the regular groups, we use vanilla SFT Howard and Ruder (2018); Raffel et al. (2023), an established and robust practice. It involves fine-tuning a pretrained LM on labeled examples using a standard supervised objective, specifically like cross-entropy. Specific prompts are provided in Table 4 in Appendix.
As to the hard group, we hypothesize that as hard questions require multiple logical steps and multiple attempts to learn core facts. So, the standard SFT is suboptimal. We propose to elicit a chain-of-thought and allow the model to incrementally build the answer step-by-step as suggested by Wei et al. (2023).
@@ 3.2 Complexity estimation approaches
To find the most suitable metric for the training pipeline, we analyze the performance of the following techniques: MASJ reasoning score, MASJ education level, a single token answer entropy, a chain-of-thought answer entropy, a chain-of-thought aggregated response entropy. Used prompts are available in Appendix A.1.
3.2.1 MASJ reasoning score
To collect the MASJ-based reasoning score, we go over the multiple choice question answering dataset and query a large auxiliary LLM for the estimate. We prompt the model to provide the number of logical steps required to answer the question: low, medium, high. Next, we query the large LLM again to estimate the quality of the previous assessment from 1 to 10 following the practice introduced in MT-Bench by Zheng et al. (2023). It allows us to filter out low quality scores by keeping only the ones rates above or equal to 9.
3.2.2 MASJ education level
As the other expert-like metrics, we ask a large LLM to estimate the required level of education to answer the question correctly. It is a natural human-like value used in other datasets Rein et al. (2023); Lu et al. (2022).
@@ 4.1 Experimental setup
Dataset
We conduct all experiments on the multiple choice question answering dataset MMLU-Pro Wang et al. (2024), widely adopted by the community as one of the main performance benchmarks. It spans 14 domains, offering a broad selection of questions of varying complexities. Each question has approximately ten options, with a single correct one, which eliminates ambiguity in evaluation.
@@
Regarding the models  [^3] All models and datasets are published under permissive licenses that allow them to be used for research purposes, we utilize a variety of open models to analyze how the trend changes with model size. At the same time, we focus on smaller models for fine-tuning to make our results reproducible and more relevant to the distillation scenario.
We apply our pipeline to two models: Qwen2.5-3B and Phi-4-Mini. For them, we measure single token and chain-of-thought entropy, collect other response metadata, fine-tune the models, and evaluate the overall pipeline.
@@ 4.1.1 Our method and baselines
In experiments, we performed training for 10 or 20 epochs, see more information below. In case of the reasoning used during training, we split training to two equal halves for the first and the second round of fine-tuning.
@@ 4.1.2 Technical details
Based on ROC AUC results and positive performance of SFT for medium and easy questions (see 4.3), we take single token entropy as our complexity metric for the pipeline described in Section 3.1.
@@ 4.2 Main results
(a) Accuracy for fine-tuning pipelines after 10 epochs for Qwen 3B
(b) Accuracy for fine-tuning pipelines after 10 epochs for Phi-4-mini
(a) Accuracy for fine-tuning pipelines after 20 epochs for Qwen 3B
(b) Accuracy for fine-tuning pipelines after 20 epochs for Phi-4-mini
Table  1 and Figures  3(a)  3(b) show the results. We see that the proposed training scheme results in a significant improvement over SFT, curriculum-based SFT and an alternative training scheme that uses distillation for only easy and medium questions. Qwen 3B achieves an accuracy of 0.50 compared to 0.47 and 0.42 for alternative and baseline, respectively, while Phi-4-mini gets to 0.60 compared to 0.58 and 0.46. At the same time, we see that the pipeline provides comparable performance to the distillation with significantly less data in number of tokens processed. Qwen 3B achieves an accuracy of $0.50$ compared to $0.49$ saving $79\%$ tokens, while Phi-4-mini gets to $0.60$ compared to $0.63$ saving $82\%$ tokens. We can also detect an uptrend in the pipeline training while the full distillation approach exhibits the signs of a plateau.
To confirm the plateau trend, we performed an extended set of experiments for 20 epochs with the same split of training epochs. Table 2 and Figures  4(a)  4(b) show the results. We find an even stronger confirmation of the previous findings with the proposed scheme outperforming all baselines. Qwen 3B achieves an accuracy of $0.52$ compared to $0.50$ and $0.39$ for distillation and baseline, respectively, while Phi-4-mini gets to $0.64$ compared to $0.63$ and $0.51$. The same data savings of $79\%$ and $82\%$ apply.
@@
Distillation 0.49 / 19.72 k 0.63 / 15.15k
Alternative 0.47 / 5.88k 0.58 / 4.91k
Pipeline (ours) 0.50 / 3.99k 0.60 / 2.67k
Table 1: Accuracy / tokens processed for complexity-aware fine-tuning pipelines after 10 epochs
Method Qwen 3B Phi4-mini
SFT 0.39 / 59k 0.51 / 54k
Distillation 0.50 / 39.45k 0.63 / 30.30k
Pipeline (ours) 0.52 / 7.98k 0.64 / 5.35k
Table 2: Accuracy / tokens processed for complexity-aware fine-tuning pipelines after 20 epochs
@@ 4.3.1 Alternative complexity estimates
To analyze the appropriate fine-tuning methods for each complexity band we start with performing the standard SFT across each group.
The same data splitting procedure as described in Section 4.2 is applied for MASJ reasoning score and single-token entropy as complexity metrics.
@@
Figures 5(d) 5(c) show the results of SFT for each group split by the single-token entropy. For Phi-4-mini, we see that medium and easy questions outperform hard ones for the first 10 epochs. Shortly after, performance starts to decline for all groups. For Qwen 3B, we do not see a significant difference between the groups. Moreover, the performance plateaus after five epochs.
Figures 5(b) and 5(a) show the results of SFT for each group split by the MASJ reasoning score. We do not see a strong difference in performance between the groups. In combination with questionable ROC AUC scores, provided in Table 9, it makes MASJ reasoning score a less favorable metric for further experiments.
@@
Figures 6(a) and 6(b) show pipeline training results with cross-entropy as a complexity metric compared to entropy. We observe comparable results for Qwen 3B, while Phi4-mini demonstrates superior performance with entropy. The difference in performance could be attributed to entropy capturing the information across the entire distribution instead of a single correct token.
@@
4.3.2 Curriculum learning fragility
Multi-stage pipelines such as curriculum learning show extreme sensitivity to hyperparameters such as a number of training epochs, learning rates and etc.
Figure 7 reveal the fragility of the original curriculum learning approach based on SFT. In this experiment we train a model on the data with increasing complexity for 10 and 20 epochs with different splits:
We see how Qwen 3B plateaus at a lower accuracy with more training epochs, while Phi4-mini quickly overfits on the data and its performance degrades.
4.3.3 Quality of complexity estimation
@@
We see there that the single token response shows the best results. In-depth analysis of ROC AUC and accuracy scores with splits across domains and larger selection of models can be found in tables 7 and 8 in the Appendix.
Key observations:
@@ 5 Conclusion and discussion
This paper introduces a complexity-aware fine-tuning pipeline that measures the model response uncertainty using the entropy of its own predicted answer and then trains it on the resulting easy, medium, and hard splits with different tactics.
Using the entropy-based data split, we find that different complexity scores require different training approaches. Standard supervised fine-tuning (SFT) is enough for the easy and medium bands, but lags on the hard band. For the hard questions, adding a distilled chain-of-thought from a large LLM unlocks further gains. Our pipeline achieves accuracies of 0.52/0.64 vs. 0.39/0.51 for Qwen 3B/Phi-4-mini, while using $81\%$ less data.
The pipeline is fully automated and can be included in other fine-tuning workflows. It suggests that curriculum ideas still matter for today’s LLMs: letting the model focus on what it can already solve directly, while giving extra guidance only where it struggles, yields a better allocation of limited model capacity.
@@
  • Proposed pipeline is tested only on MMLU-Pro and small models. Results may change for other question answering datasets, open-ended tasks, other domains, or larger LLMs.
@@
  • We split data into 3 equal parts and did not explore other possible boundaries.
  • We did not conduct an extensive ablation study which might reveal that our approach does not suggest the best possible combination or sequence of training within the current framework. It remains an area for further research.

### v3

@@ Abstract
General-purpose Large Language Models (LLMs) are frequently fine-tuned through supervised fine-tuning (SFT) to enhance performance in specific domains. Better results can be achieved by distilling the chain-of-thought of a larger model at the cost of numerous expensive calls and a much greater amount of data. We propose a novel blueprint for efficient fine-tuning that uses reasoning only for complex data identified by entropy. Specifically, across three small open models ($\approx 3B$) we split the training data into complexity categories by a single token answer entropy (ROC AUC $0.73$), fine-tune large language models (LLMs) via SFT and distillation, and show that our pipeline significantly outperforms the standard SFT approach ($0.58$ vs $0.45$ average accuracy) and outperforms the distillation approach ($0.58$ vs $0.56$ average accuracy) while using $81\%$ less data. We publish our code [^1] https://github.com/LabARSS/complexity-aware-fine-tuning and data [^2] https://github.com/LabARSS/complexity-aware-fine-tuning?tab=readme-ov-file#data to facilitate further research in this direction.
@@ 1 Introduction
To address this gap, we propose a fully automated pipeline, presented in Figure 1, that focus on both accuracy and efficiency of the fine-tuning procedure. It consists of two steps: (1) a split of the data for fine-tuning by the complexity of the task to regular and hard questions and (2) training a model via supervised fine-tuning (SFT) followed by another round of SFT that utilizes distillation on a chain-of-thought from a larger model Li et al. (2023); Hsieh et al. (2023). We confirm the effectiveness of the framework by fine-tuning three open models: Qwen2.5-3B Yang et al. (2024a), Phi-4-Mini Microsoft et al. (2025), LLama 3.2 3B Grattafiori et al. (2024), on three datasets: multiple choice question answering dataset MMLU-Pro, multiple choice question answering dataset MedMCQA, mathematical dataset GSM8K.
@@
  • A training pipeline to fine-tune LLMs via SFT and distillation for regular and hard data correspondingly. Our procedure provides accuracies:
    • (MMLU-Pro) $0.52$ and $0.64$ c.t. $0.39$ and $0.51$ for SFT and $0.50$ and $0.63$ for distillation baselines.
    • (GSM8K) $0.82$ and $0.89$ c.t. $0.13$ and $0.19$ for SFT and $0.79$ and $0.89$ for distillation baselines.
    • (MedMCQA) $0.56$ and $0.58$ c.t. $0.51$ and $0.51$ for SFT and $0.56$ and $0.57$ for distillation baselines.
@@ 3.1 Training pipeline
2. Data aggregation.
To aggregate the data into groups by complexity (regular, hard), we evenly divide the dataset into three parts, ordering the entries by entropy of the response. A group with lower entropy values is categorized as easy and with the higher values as hard.
3. Fine-tuning.
We propose applying different fine-tuning strategies based on the complexity of the data.
For the regular groups, we use vanilla SFT Howard and Ruder (2018); Raffel et al. (2023), an established and robust practice. It involves fine-tuning a pretrained LM on labeled examples using a standard supervised objective, specifically, cross-entropy. Used prompts are provided in Table 4 in Appendix.
As to the hard group, we hypothesize that hard questions require multiple logical steps and multiple attempts to learn core facts. So, the standard SFT is suboptimal. We propose to elicit a chain-of-thought and allow the model to incrementally build the answer step-by-step as suggested by Wei et al. (2023).
@@
Pseudo-code for the algorithm is provided below in Algorithm 1.
Algorithm 1 Complexity-Aware Fine-Tuning Pipeline
1:Training dataset $\mathcal{D}=\{(q_{i},a_{i})\}_{i=1}^{N}$, Student LLM $\mathcal{M}_{S}$, Teacher LLM $\mathcal{M}_{T}$
2:Fine-tuned Student LLM $\mathcal{M}_{S}^{*}$
3:// Step 1: Uncertainty Estimation
4:for each $(q_{i},a_{i})\in\mathcal{D}$ do
5:  $\mathbf{p}_{i}\leftarrow\mathcal{M}_{S}(q_{i})$ $\triangleright$ Get token probability distribution
6:  $h_{i}\leftarrow-\sum_{j=1}^{|\mathcal{V}|}p_{ij}\log p_{ij}$ $\triangleright$ Compute answer token entropy
7:end for
8:
9:// Step 2: Data Aggregation by Complexity
10:Sort $\mathcal{D}$ by entropy values $\{h_{i}\}_{i=1}^{N}$ in ascending order
[…]
15:
16:// Step 3: Standard SFT on Regular Data
17:for epoch =1=1 to $E_{1}$ do
18:  for batch $\mathcal{B}\subset\mathcal{D}_{\text{regular}}$ do
19:   $\mathcal{L}_{\text{SFT}}\leftarrow-\sum_{(q,a)\in\mathcal{B}}\log P_{\mathcal{M}_{S}}(a\mid q)$
20:   $\mathcal{M}_{S}\leftarrow\mathcal{M}_{S}-\eta\nabla\mathcal{L}_{\text{SFT}}$
21:  end for
22:end for
23:
24:// Step 4: Distill Chain-of-Thought for Hard Data
25:for each $(q_{i},a_{i})\in\mathcal{D}_{\text{hard}}$ do
26:  $r_{i}\leftarrow\mathcal{M}_{T}(q_{i})$ $\triangleright$ Generate CoT reasoning from teacher
27:  $\mathcal{D}_{\text{hard}}^{\text{CoT}}\leftarrow\mathcal{D}_{\text{hard}}^{\text{CoT}}\cup\{(q_{i},r_{i},a_{i})\}$
28:end for
29:
30:// Step 5: SFT with Reasoning on Hard Data
31:for epoch =1=1 to $E_{2}$ do
32:  for batch $\mathcal{B}\subset\mathcal{D}_{\text{hard}}^{\text{CoT}}$ do
33:   $\mathcal{L}_{\text{CoT}}\leftarrow-\sum_{(q,r,a)\in\mathcal{B}}\log P_{\mathcal{M}_{S}}(r\oplus a\mid q)$
34:   $\mathcal{M}_{S}\leftarrow\mathcal{M}_{S}-\eta\nabla\mathcal{L}_{\text{CoT}}$
35:  end for
36:end for
37:return $\mathcal{M}_{S}^{*}\leftarrow\mathcal{M}_{S}$
@@ 3.2 Complexity estimation approaches
To find the most suitable metric for the training pipeline, we analyze the performance of the following techniques: a single token answer entropy, a chain-of-thought answer entropy, a chain-of-thought aggregated response entropy, MASJ reasoning score, and MASJ education level. Used prompts are available in Appendix A.1.
@@ 3.2.1 Answer entropy
3.2.3 MASJ reasoning score
To collect the MASJ-based reasoning score, we go over the multiple-choice question answering dataset and query a large auxiliary LLM for the estimate. We prompt the model to provide the number of logical steps required to answer the question: low, medium, or high. Next, we query the large LLM again to estimate the quality of the previous assessment from 1 to 10, following the practice introduced in MT-Bench by Zheng et al. (2023). It allows us to filter out low quality scores by keeping only the ones with rates above or equal to 9.
3.2.4 MASJ education level
Like the other expert-like metrics, we ask a large LLM to estimate the required level of education to answer the question correctly. It is a natural human-like value used in other datasets Rein et al. (2023); Lu et al. (2022).
@@ 4.1 Experimental setup
Datasets
We conduct experiments across multiple datasets:
  • Multiple choice question answering dataset MMLU-Pro Wang et al. (2024), widely adopted by the community as one of the main performance benchmarks. It spans 14 domains, offering a broad selection of questions of varying complexities. Each question has approximately ten options, with a single correct one, which eliminates ambiguity in evaluation.
  • GSM8K (Grade School Math 8K) Cobbe et al. (2021) - dataset of 8.5K high quality linguistically diverse grade school math word problems.
  • MedMCQA Pal et al. (2022) - large-scale, Multiple-Choice Question Answering (MCQA) dataset designed to address real-world medical entrance exam questions. It has more than 194k high-quality AIIMS and NEET PG entrance exam MCQs covering 2.4k healthcare topics and 21 medical subjects are collected with an average token length of 12.77 and high topical diversity
@@
Regarding the models  [^3] All models and datasets are published under permissive licenses that allow them to be used for research purposes, we utilize a variety of open models to analyze how the trend changes with model size. At the same time, our focus is on smaller models for fine-tuning to make our results reproducible and more relevant to the distillation scenario.
We apply our pipeline to three models: Qwen2.5-3B, Phi-4-Mini, LLama 3.2 3B. For them, we measure single token and chain-of-thought entropy, collect other response metadata, fine-tune the models, and evaluate the overall pipeline.
@@ 4.1.1 Our method and baselines
In experiments, we performed training for 10 or 20 epochs; see more information below. In the case of the reasoning used during training, we split the training into two equal halves for the first and the second round of fine-tuning.
@@ 4.1.2 Technical details
Based on ROC AUC results and SFT’s positive performance on medium and easy questions (see 4.3), we use single-token entropy as our complexity metric for the pipeline described in Section 3.1.
@@ 4.2 Main results
(a) Accuracy for fine-tuning pipelines after 10 epochs for Qwen 3B (MMLU-Pro)
(b) Accuracy for fine-tuning pipelines after 10 epochs for Phi-4-mini (MMLU-Pro)
(a) Accuracy for fine-tuning pipelines after 20 epochs for Qwen 3B (MMLU-Pro)
(b) Accuracy for fine-tuning pipelines after 20 epochs for Phi-4-mini (MMLU-Pro)
Table  1 and Figures  2(a)  2(b) show the results for MMLU Pro. We see that the proposed training scheme results in a significant improvement over SFT, curriculum-based SFT, and an alternative training scheme that uses distillation for only easy and medium questions. Qwen 3B achieves an accuracy of 0.50 compared to 0.47 and 0.42 for alternative and baseline, respectively, while Phi-4-mini gets to 0.60 compared to 0.58 and 0.46. At the same time, we see that the pipeline provides comparable performance to the distillation with significantly less data in the number of tokens processed. Qwen 3B achieves an accuracy of $0.50$ compared to $0.49$, saving $79\%$ tokens, while Phi-4-mini gets to $0.60$ compared to $0.63$, saving $82\%$ tokens. We can also detect an uptrend in the pipeline training, while the full distillation approach exhibits the signs of a plateau.
To confirm the plateau trend, we ran an extended set of experiments over 20 epochs with the same training split. Table 2 and Figures  3(a)  3(b) show the results. We find an even stronger confirmation of the previous findings with the proposed scheme outperforming all baselines. Qwen 3B achieves an accuracy of $0.52$ compared to $0.50$ and $0.39$ for distillation and baseline, respectively, while Phi-4-mini gets to $0.64$ compared to $0.63$ and $0.51$. The same data savings of $79\%$ and $82\%$ apply.
To confirm that the method generalizes to other datasets and models, we perform the same experiments for GSM8K and MedMCQA datasets across two models: LLama 3.2 3B and Phi-4 mini. Table 2 shows the results. We see that across both models and datasets, our pipeline achieves accuracy either comparable to or surpassing the distillation baseline, with significant data savings.
Training hyperparameters are provided in the Appendix D.
@@
Distillation 0.49 / 19,72k 0.63 / 15,15k
Alternative 0.47 / 58k 0.58 / 49k
Pipeline (ours) 0.50 / 3,99k 0.60 / 2,68k
Table 1: Accuracy / tokens processed for complexity-aware fine-tuning pipelines after 10 epochs (MMLU Pro)
MMLU Pro GSM8K MedMCQA
Method Qwen 3B Phi4-mini LLama 3B Phi4-mini LLama 3B Phi4-mini
SFT 0.39 0.51 0.13 0.19 0.51 0.51
Distillation 0.50 0.63 0.79 0.89 0.56 0.57
Pipeline (ours) 0.52 0.64 0.82 0.89 0.56 0.58
Table 2: Accuracy for complexity-aware fine-tuning pipelines after 20 epochs.
@@ 4.3.1 Alternative complexity estimates
To analyze the appropriate fine-tuning methods for each complexity band, we first perform the standard SFT for each group.
The same data-splitting procedure described in Section 4.2 is applied to MASJ reasoning score and single-token entropy as complexity metrics.
@@
Figures 4(d) 4(c) show the results of SFT for each group split by the single-token entropy. For Phi-4-mini, we see that medium and easy questions outperform hard ones for the first 10 epochs. Shortly after, performance starts to decline for all groups. For Qwen 3B, we do not see a significant difference between the groups. Moreover, the performance plateaus after five epochs.
Figures 4(b) and 4(a) show the results of SFT for each group split by the MASJ reasoning score. We do not see a strong difference in performance between the groups. In combination with questionable ROC AUC scores, provided in Table 9, it makes MASJ reasoning score a less favorable metric for further experiments.
@@
Figures 5(a) and 5(b) show pipeline training results with cross-entropy as a complexity metric compared to entropy. We observe comparable results for Qwen 3B, while Phi4-mini demonstrates superior performance with entropy. The difference in performance could be attributed to entropy capturing the information across the entire distribution instead of a single correct token.
@@
4.3.2 Curriculum learning fragility
Multi-stage pipelines such as curriculum learning show extreme sensitivity to hyperparameters such as the number of training epochs, learning rates, etc.
Figure 6 reveals the fragility of the original curriculum learning approach based on SFT. In this experiment, we train a model on the data with increasing complexity for 10 and 20 epochs with different splits:
Qwen 3B plateaus at a lower accuracy with more training epochs, while Phi4-mini quickly overfits on the data, and its performance degrades.
4.3.3 Quality of complexity estimation
@@
We see that the single-token response yields the best results. In-depth analysis of ROC AUC and accuracy scores with splits across domains and a larger selection of models can be found in tables 7 and 8 in the Appendix.
Key observations include:
@@
5 Acknowledgments
The research was supported by the Russian Science Foundation grant No. 25-11-00355, https://rscf.ru/project/25-11-00355/.
@@ 6 Conclusion and discussion
This paper introduces a complexity-aware fine-tuning pipeline that measures model response uncertainty using the entropy of its predicted answer, then trains on the resulting easy, medium, and hard splits via different tactics.
@@
Using the entropy-based data split, we find that different complexity scores require different training approaches. Standard supervised fine-tuning (SFT) is enough for the easy and medium bands, but lags on the hard band. For hard questions, adding a distilled chain of thought from a large LLM unlocks further gains. Our pipeline achieves accuracies of 0.52/0.64 vs. 0.39/0.51 for Qwen 3B/Phi-4-mini, while using $81\%$ less data.
The pipeline is fully automated and can be included in other fine-tuning workflows. Our findings suggest that curriculum ideas still matter for today’s LLMs: letting the model focus on what it can already solve directly, while giving extra guidance only where it struggles, yields a better allocation of limited model capacity.
@@
  • The proposed pipeline is tested only on small models. Results may differ across other question-answering datasets, open-ended tasks, other domains, or larger LLMs.
@@
  • We split the data into 3 equal parts and did not explore other possible boundaries.
  • We did not conduct an extensive ablation study, which might reveal that our approach does not yield the best possible training combination or sequence within the current framework. It remains an area for further research.